OpenStudy (anonymous):

Compare the data in the table and the equation. x y 4 12 5 15 6 18 Equation: y = 3x Which representation has the greatest slope? The equation has the greatest slope. The table has the greatest slope. The table and equation have the same slope. Their slopes cannot be determined.

OpenStudy (anonymous):

@nincompoop can you please help me?

OpenStudy (anonymous):

@jabez177 pls help me

OpenStudy (anonymous):

@Compassionate can you please help me?

OpenStudy (anonymous):

@AlexandervonHumboldt2 can you please help me?

OpenStudy (anonymous):

Please I really need help

OpenStudy (dayakar):

slope=y/x compare it u can get the answer

OpenStudy (anonymous):

I don't know what the question means by compare it, thats what I'm confused about @dayakar

OpenStudy (dayakar):

i explain first one m=y/x = 12/4 =3

OpenStudy (anonymous):

oh ok that make so much more sense